# ArtNet: Hierarchical Clustering-Based Artificial Netlist Generator for ML and DTCO Applications

### v1

@@ IV-E Assessment in DTCO Context: Mini-Brains
We explore a novel mini-brains approach [15] that enables more efficient design-technology co-optimization, helping to mitigate the increasing complexity and scale of real modern designs. Our mini-brain methodology generates artificial netlists scaled to 10% of a full-scale design’s size, with the goal of replicating the PPA characteristics of full-scale designs while significantly reducing P&R runtimes.
@@ IV-E1 PPA-Matching Evaluation
TABLE XVI: Loss and runtime reduction results.
  Design MAPE Loss
Runtime
Reduction
jpeg 2.06% -79.26%
nova 7.55% -91.45%
netcard 3.74% -93.74%
tate_pairing 3.43% -88.58%
CA53 4.70% -87.50%
ariane-133 3.50% -95.53%
To assess reliability, we conduct robustness tests by sweeping target clock periods and calculating the average of two metrics; the MAPE of the effective clock period and the MAPE of total power per area. The overall loss is defined as:
@@
$\text{MAPE}_{\text{power}}=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{P_{\text{real}}^{(i)}-P_{\text{mini}}^{(i)}}{P_{\text{thumb}}^{(i)}}\right|\times 100.$ (7)
Here, $T_{\text{real}}^{(i)}$ and $T_{\text{mini}}^{(i)}$ respectively denote the effective clock periods of the real design and its mini-brain for the $i^{th}$ sample. Likewise, $P_{\text{real}}^{(i)}$ and $P_{\text{mini}}^{(i)}$ respectively denote the total power per area of the real design and its mini-brain. Table XVI shows the potential of mini-brains to accurately replicate PPA metrics while significantly reducing runtime compared to full-scale designs, particularly in the DTCO context.
@@ IV-E2 BEOL Parameter Evaluation
Ranking Consistency Evaluation. We evaluate whether mini-brains can reliably preserve the relative ranking of design outcomes under different BEOL settings. For each mini-brain and each real design, we rank all 81 BEOL configurations in ascending order of total power. For each full design and corresponding mini-brain, we evaluate rank correlation of outcomes using Kendall and Spearman correlation coefficients. Table XVIII shows consistently high rank correlation across diverse designs, indicating that mini-brains successfully capture the sensitivity of design quality to BEOL variation. This confirms that mini-brains are not only efficient to generate but also accurate enough to guide BEOL parameter exploration.
TABLE XVIII: DTCO mini-brain evaluation results. Design quality rank correlation between real and artificial designs, across 81 BEOL parameter combinations.
  Design Kendall Spearman
jpeg 0.621 0.823
nova 0.662 0.856
netcard 0.623 0.802
tate_pairing 0.760 0.931
CA53 0.752 0.915
ariane-133 0.646 0.851

### v2

@@ IV-E Assessment in DTCO Context: Mini-Brains
We explore a novel mini-brains approach [17] that enables more efficient design-technology co-optimization, helping to mitigate the increasing complexity and scale of real modern designs. Our mini-brain methodology generates artificial netlists scaled to 10% of a full-scale design’s size, by reducing only circuit-size–related parameters ($N_{insts}$, $N_{PI}$, $N_{PO}$ and $N_{macro}$), with the goal of replicating the PPA characteristics of full-scale designs while significantly reducing P&R runtimes. The reduction in implementation flow runtime is primarily due to the reduced netlist size.
@@ IV-E1 PPA-Matching Evaluation
TABLE XVI: MAPE loss and runtime reduction results.
  Design MAPE Loss (%) P&R Runtime Reduction
ANG GNL ArtNet
jpeg 2.66 18.17 2.06 -79.26%
nova 27.55 133.37 7.33 -91.45%
netcard 31.37 14.19 3.74 -93.74%
tate_pairing 11.58 34.44 3.43 -88.58%
CA53 N/A 8.26 4.70 -87.50%
ariane-133 N/A 169.61 3.50 -95.53%
To assess reliability, we conduct robustness tests by sweeping target clock periods and calculating the average of two metrics; the MAPE of the effective clock period and the MAPE of total power per area. Fig. 12 shows that the mini-brain closely tracks the PPA trends of the real design across a wide range of target clock periods. On the y-axis, power values are divided by area to enable comparison between designs of different sizes. The overall loss is defined as:
@@
$\text{MAPE}_{\text{power}}=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{P_{\text{real}}^{(i)}-P_{\text{mini}}^{(i)}}{P_{\text{real}}^{(i)}}\right|\times 100.$ (7)
Here, $T_{\text{real}}^{(i)}$ and $T_{\text{mini}}^{(i)}$ respectively denote the effective clock periods of the real design and its mini-brain for the $i^{th}$ sample. Likewise, $P_{\text{real}}^{(i)}$ and $P_{\text{mini}}^{(i)}$ respectively denote the total power per area of the real design and its mini-brain. Table XVI shows the potential of mini-brains to accurately replicate PPA metrics while significantly reducing P&R runtime compared to full-scale designs, particularly in the DTCO context.
@@ IV-E2 BEOL Parameter Evaluation
Ranking Consistency Evaluation. We evaluate whether mini-brains can reliably preserve the relative ranking of design outcomes under different BEOL settings. For each mini-brain and each real design, we rank all 81 BEOL configurations in ascending order of total power. For each full design and corresponding mini-brain, we evaluate rank correlation of outcomes using Kendall and Spearman correlation coefficients. Table XVIII shows consistently high rank correlation and low MAPE for power per unit area across diverse designs, indicating that mini-brains successfully capture the sensitivity of design quality to BEOL variation. This confirms that mini-brains are not only efficient to generate but also accurate enough to guide BEOL parameter exploration.
TABLE XVIII: DTCO mini-brain evaluation results. Design quality rank correlation between real and artificial designs, across 81 BEOL parameter combinations and MAPE for power per unit area ($\muW/μ\mum$2).
  Design Kendall Spearman MAPE (%)
jpeg 0.621 0.823 2.37
nova 0.662 0.856 5.21
netcard 0.623 0.802 1.14
tate_pairing 0.760 0.931 4.98
CA53 0.752 0.915 7.26
ariane-133 0.646 0.851 2.99